# Providing Safety Assurances for Systems with Unknown Dynamics

### v2

@@ Abstract
I Introduction
@@
Our method is shown to provide significant benefits as it explicitly incorporates state and control-dependent model uncertainty in safety analysis and, as a result, provides more robust, but not overly conservative controllers for the system. To summarize, the key contributions of this letter are two-fold: 1) We propose a framework for providing robust safety assurances to systems with unknown dynamics by incorporating state and control-dependent model uncertainty in HJ reachability analysis. We also provide a closed-form solution to the two-player zero-sum differential game between the control and disturbance player, in order to accommodate the state and control dependent nature of the model uncertainty, and 2) We provide a concrete instantiation of our framework by modeling the system with an ensemble of neural networks and solving a robust optimal control problem to safeguard the learned model against bounded model uncertainty.
II Problem Formulation
@@
where $x\in\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ and $u\in\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$ are the state and control of the system, and $f_{1}(x)$ and $f_{2}(x)$ are matrices of appropriate sizes. Given a set of states $\mathcal{F}\subset\mathcal{X}$, which we refer to as the failure set, that the system must avoid (e.g., obstacles for a mobile robot), our goal is ensuring the system $f$ does not enter the failure set $\mathcal{F}$ for predefined time horizon $[0,T]$. More formally as state in Prob. 2.
Problem 1 (Safety Problem for System $f$).
@@
Problem 2 (Safety Assurance under Model Uncertainty).
Given a failure set $\mathcal{F}$ and time horizon $[0,T]$, obtain a set $\mathcal{R}\subseteq\mathcal{X}$, along with a state-feedback controller $\pi_{\mathcal{R}}$, such that under the condition that all realized model uncertainties are contained within $\mathcal{D}$, system $f$ starts within $\hat{\mathcal{S}}$ implies that $f$ does not enter $\mathcal{F}$ for $[0,T]$ using controls from $\hat{\pi}$, or equivalently, $\mathcal{R}\subseteq\mathcal{S}^{*}$.
Definition 1.
@@
III Background
In this section, we provide a brief overview of Hamtilon-Jacobi (HJ) reachability analysis, an approach that can help solve Prob. 2. Let $g$ be a system described by dynamics $\dot{x}=g(x,u,d)$, where $x\in\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$, $u\in\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$, and $d\in\mathcal{D}\subseteq\mathbb{R}^{n_{d}}$ are the state, control, and disturbance of the system. Disturbance $d$, in the interest of this letter, describes the model uncertainty.We use $\mathbf{u}:[0,T]\rightarrow\mathcal{U}$ and $\mathbf{d}:[0,T]\rightarrow\mathbb{R}^{n_{d}}$ to denote the control and disturbance signals. Furthermore, we denote the state trajectory starting from state $x$ at time $t$ evolved with control and disturbance signals $\mathbf{u}(\cdot)$ and $\mathbf{d}(\cdot)$ as $\xi_{x,t}^{\mathbf{u},\mathbf{d}}$. With a slight abuse of the notation, we use $\xi_{x,t}^{\mathbf{u},\mathbf{d}}(\tau)$ to denote the state at time $\tau\geq t$ along the trajectory $\xi_{x,t}^{\mathbf{u},\mathbf{d}}$.
Suppose $g(x,u,d)$ is uniformly continuous, bounded, and Lipschitz continuous in $x$ for fixed $u$ and $d$. We further assume that $\mathcal{U}$ and $\mathcal{D}$ are compact, and $\mathbf{u}(\cdot)$ as well as $\mathbf{d}(\cdot)$ are measurable. Let $l(x)$ be the signed distance function to $\mathcal{F}$. We can obtain $\mathcal{R}^{*}$ by solving the following robust optimal control problem (Prob. 3) with initial condition $x=x_{0}\ \forall x_{0}\in\mathcal{X}$ and $t=0$. In robust control literature, Prob. 3 is posed as a two-player zero-sum differential game between the control $u$ and disturbance $d$, who uses only nonanticipative strategies [18]. Let $\Gamma(t)$ be the set of nonanticipative strategies, and $\mathbb{U}(t)$ be the set of admisable control signals.
Problem 3.
@@ III Background
HJ reachability analysis provides a tractable means to compute the value function $V(x,t)$. It has been shown that $V(x,t)$ is the unique viscosity solution of the Hamilton-Jacobi-Issacs Variational Inequality (HJI-VI) [17, 18]:
@@
${\color[rgb]{0,0,0}\pi_{\mathcal{R}^{*}}(x,t)=\arg\max_{u\in\mathcal{U}}\min_{d\in\mathcal{D}}\langle\nabla V(x,t),g(x,u,d)\rangle}$ (5)
IV Safety Assurances for Learned Dynamics
@@
IV-A Model Representation and Hamiltonian Formulation
@@
IV-B Ensemble Dynamics Representation
@@
V Experiments
V-A Inverted Pendulum
@@
For the number of training samples $M=300$, we visualize the recovered safe sets for our method, all the baselines, and the ground truth in Fig. 1. The ground truth safe set, computed using the analytical expression of the dynamics, is shaded in green. The safe set recovered using our method is entirely contained within the ground truth safe set, indicating satisfaction of the safety constraint. On the other hand, Baseline 1 fails to satisfy the safety constraint, since the recovered safe set is not contained within the ground truth safe set. The safe set from Baseline 2 is not visualized in Fig. 1 because its recovered safe set is empty due to model uncertaintity bounds $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ being too conservative. Specifically, conformal prediction provides state-independent uncertainty bounds, which are dictated by the worst-case modeling errors across all states, leading to overly conservative behaviors. Baseline 3 is less conservative than Baseline 2, as its model uncertainty bounds are state-dependent, but it is more conservative than our methods since its model uncertainty $d_{3}(x)$ is control-independent and its bound $\mathcal{D}_{3}(x)$ is an overapproximation of that of our method.
@@
V-B Turtlebot Hardware Experiment

### v3

@@ Abstract
{IEEEkeywords}
Autonomous systems, Robust control, Uncertain systems
1 Introduction
[…]
@@ 1 Introduction
Our method is shown to provide significant benefits as it explicitly incorporates state and control-dependent model uncertainty in safety analysis and, as a result, provides more robust, but not overly conservative controllers for the system. To summarize, the key contributions of this letter are two-fold: 1) We propose a framework for providing robust safety assurances to systems with unknown dynamics by incorporating state and control-dependent model uncertainty in HJ reachability analysis. We also provide a closed-form solution to the two-player zero-sum differential game between the control and disturbance player, in order to accommodate the state and control-dependent nature of the model uncertainty, and 2) We provide a concrete instantiation of our framework by modeling the system with an ensemble of neural networks and solving a robust optimal control problem to safeguard the learned model against bounded model uncertainty.
2 Problem Formulation
@@
where $x\in\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$ and $u\in\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$ are the state and control of the system, and $f_{1}(x)$ and $f_{2}(x)$ are matrices of appropriate sizes. Given a set of states $\mathcal{F}\subset\mathcal{X}$, which we refer to as the failure set, that the system must avoid (e.g., obstacles for a mobile robot), our goal is ensuring the system $f$ does not enter the failure set $\mathcal{F}$ for predefined time horizon $[0,T]$. More formally as stated in Prob. 1.
Problem 1 (Safety Problem for System $f$)
@@
Problem 2 (Safety Assurance under Model Uncertainty)
Given a failure set $\mathcal{F}$ and time horizon $[0,T]$, obtain a set $\mathcal{R}\subseteq\mathcal{X}$, along with a state-feedback controller $\pi_{\mathcal{R}}$, such that under the condition that all realized model uncertainties are contained within $\mathcal{D}$, system $f$ starts within $\mathcal{R}$ implies that $f$ does not enter $\mathcal{F}$ for $[0,T]$ using controls from $\pi_{\mathcal{R}}$, or equivalently, $\mathcal{R}\subseteq\mathcal{S}^{*}$.
Definition 1
@@
3 Background
In this section, we provide a brief overview of Hamilton-Jacobi (HJ) reachability analysis, an approach that can help solve Prob. 2. Let $g$ be a system described by dynamics $\dot{x}=g(x,u,d)$, where $x\in\mathcal{X}\subseteq\mathbb{R}^{n_{x}}$, $u\in\mathcal{U}\subseteq\mathbb{R}^{n_{u}}$, and $d\in\mathcal{D}\subseteq\mathbb{R}^{n_{d}}$ are the state, control, and disturbance of the system. Disturbance $d$, in the interest of this letter, describes the model uncertainty. We use $\mathbf{u}:[0,T]\rightarrow\mathcal{U}$ and $\mathbf{d}:[0,T]\rightarrow\mathbb{R}^{n_{d}}$ to denote the control and disturbance signals. Furthermore, we denote the state trajectory starting from state $x$ at time $t$ evolved with control and disturbance signals $\mathbf{u}(\cdot)$ and $\mathbf{d}(\cdot)$ as $\xi_{x,t}^{\mathbf{u},\mathbf{d}}$. With a slight abuse of the notation, we use $\xi_{x,t}^{\mathbf{u},\mathbf{d}}(\tau)$ to denote the state at time $\tau\geq t$ along the trajectory $\xi_{x,t}^{\mathbf{u},\mathbf{d}}$.
Suppose $g(x,u,d)$ is uniformly continuous, bounded, and Lipschitz continuous in $x$ for fixed $u$ and $d$. We further assume that $\mathcal{U}$ and $\mathcal{D}$ are compact, and $\mathbf{u}(\cdot)$ as well as $\mathbf{d}(\cdot)$ are measurable. Let $l(x)$ be the signed distance function to $\mathcal{F}$. We can obtain $\mathcal{R}^{*}$ by solving the following robust optimal control problem (Prob. 3) with initial condition $x=x_{0}\ \forall x_{0}\in\mathcal{X}$ and $t=0$. In robust control literature, Prob. 3 is posed as a two-player zero-sum differential game between the control $u$ and disturbance $d$, who uses only nonanticipative strategies [18]. Let $\Gamma(t)$ be the set of nonanticipative strategies, and $\mathbb{U}(t)$ be the set of admissible control signals.
Problem 3 (Robust Safety Optimal Control Problem)
@@ 3 Background
HJ reachability analysis provides a tractable means to compute the value function $V(x,t)$. It has been shown that $V(x,t)$ is the unique viscosity solution of the Hamilton-Jacobi-Isaacs Variational Inequality (HJI-VI) [17, 18]:
@@
$\pi_{\mathcal{R}^{*}}(x,t)=\arg\max_{u\in\mathcal{U}}\min_{d\in\mathcal{D}}\langle\nabla V(x,t),g(x,u,d)\rangle$ (5)
4 Safety Assurances for Learned Dynamics
@@
4.1 Model Representation and Hamiltonian Formulation
@@
4.2 Ensemble Dynamics Representation
@@
5 Experiments
5.1 Inverted Pendulum
@@
For the number of training samples $M=300$, we visualize the recovered safe sets for our method, all the baselines, and the ground truth in Fig. 1. The ground truth safe set, computed using the analytical expression of the dynamics, is shaded in green. The safe set recovered using our method is entirely contained within the ground truth safe set, indicating satisfaction of the safety constraint. On the other hand, Baseline 1 fails to satisfy the safety constraint, since the recovered safe set is not contained within the ground truth safe set. The safe set from Baseline 2 is not visualized in Fig. 1 because its recovered safe set is empty due to model uncertainty bounds $\mathcal{D}_{1}$ and $\mathcal{D}_{2}$ being too conservative. Specifically, conformal prediction provides state-independent uncertainty bounds, which are dictated by the worst-case modeling errors across all states, leading to overly conservative behaviors. Baseline 3 is less conservative than Baseline 2, as its model uncertainty bounds are state-dependent, but it is more conservative than our methods since its model uncertainty $d_{3}(x)$ is control-independent and its bound $\mathcal{D}_{3}(x)$ is an overapproximation of that of our method.
@@
5.2 Turtlebot Hardware Experiment